A particle moves with constant acceleration. Let V1,V2,V3 be the average velocities in the successive intervals of time t1,t2,t3. then:
V1-V2=? And V2-V3=?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Assume the particles moves the separation AB, BC and CD in time t1,t2 and t3 individually. 
Speed at B = u+at1 
Speed at C = u+a(t1+t2) 
Speed at D = u + a( t1+t2+t3) 
Normal speed (v1)=( starting +final)/2 
= (u + u + at1)/2 
= u+1/2 at1 
V2 = u+at1+1/2at2 
V3 = u+ at1+at2+1/2at3 
There fore we get 
(V1-V2)/(V2-V3) = (t1+t2)/(t2+t3)
 
 
Regards
Arun (askIITians forum expert)